Calculate the change in length of a 100mm aluminium bar that has increased in temperature by 55⁰C.  

Take the coefficient of expansion to be 25 x 10-6k-1

1
Expert's answer
2022-02-20T08:20:55-0500

Given:

"l_0=100\;\rm mm"

"\Delta T=55\:\rm K"

"\alpha=25*10^{-6}\:\rm K^{-1}"


The thermal linear expansion is described by an equation


"\Delta l=l_0\alpha\Delta T\\
={\rm100\: mm*25*10^{-6}\:\rm K^{-1}*55\: K}\\
=\rm0.014\: mm"